OpenStudy (anonymous):

y varies directly with x. When y = 12, x = 16. Find x when y = 24.

OpenStudy (lalaly):

y=kx when y=12 x=16 12=k*16 k=12/16 = 3/4 when y=24 24=3/4 *x x=24*4 / 3

OpenStudy (anonymous):

32 18 64

OpenStudy (lalaly):

32

OpenStudy (anonymous):

thanksss

OpenStudy (lalaly):

np:D